# Second- and third-order properties of multidimensional Langevin equations

### v14

@@ 6.1. Preliminaries
showing that we only need the coefficient of $x^{3}$ to be $\mathcal{O}(a^{2})$ for the model to make sense, a result which could have been anticipated by symmetry. Thus, for small $a$, we can interpret the quadratic model for $\langle\dot{x}\mid x\rangle$ irrespective of any higher-order contributions. We also assume the linear model for inhomogeneous diffusion, noting similarly that in order for the diffusion function to be non-negative we can add a term $(b^{2}/4)x^{2}$, which is $\mathcal{O}(b^{2})$. Thus, to order $\mathcal{O}((a,b)^{2})$, our model is:
@@ 6.2. One-dimensional case
$\qquad{}+a^{2}\mathcal{O}(1){He}_{3}(x)+a^{3}\mathcal{O}(1){He}_{4}(x)+\cdots,$
@@
$\qquad{}+(2a+a^{2}\mathcal{O}(a,b)){He}_{3}(x)+a^{2}\mathcal{O}(1){He}_{4}(x)+\cdots,$
@@ 8.3.2. Solution of third-order quantities
$\quad{}+\mathcal{O}(\mu^{-1})\mathcal{O}(a_{vv},b_{v})+\mathcal{O}(\mu^{-2})\mathcal{O}(a_{xx},a_{xv},b_{x}),$
@@ 8.3.3. Quantitative significance
which is also a reasonable choice. Lastly, it should be noted that $\langle x^{i}v^{j}v^{k}\rangle=L(x^{i}v^{j},x^{k})/2$ and $\langle v^{i}v^{j}v^{k}\rangle=L(v^{i}v^{j},x^{k})/2$ are also angular momenta, but the calculated ensemble covariances using the angular momenta vanish when $\mathbf{x}$ and $\mathbf{v}$ are expressed in terms of $\widetilde{\mathbf{x}}$ and $\widetilde{\mathbf{v}}$ and when Eqs. (361)–(362) are applied.
@@ 8.4. Dynamics on long time-scales
On $\mathcal{O}(\mu)$ time-scales, the process is governed by the dynamics of $\widetilde{\mathbf{x}}$, with relative error $\mathcal{O}(\mu^{-2})$. For example, if $\tau\sim\mathcal{O}(\mu)$, then:
(366) $\frac{\langle x^{i}(0)x^{j}(0)x^{k}(\tau)-x^{i}(\tau)x^{j}(\tau)x^{k}(0)\rangle}{\tau}=\frac{\langle\widetilde{x}^{i}(0)\widetilde{x}^{j}(0)\widetilde{x}^{k}(\tau)-\widetilde{x}^{i}(\tau)\widetilde{x}^{j}(\tau)\widetilde{x}^{k}(0)\rangle}{\tau}+\mathcal{O}(\mu^{-3}).$
We now show that the dynamics of $\widetilde{\mathbf{x}}$ is Markovian on these time-scales, with relative error $\mathcal{O}(\mu^{-2})$.
We make the assumption $\left|(a,b)\right|\lesssim\mathcal{O}(\mu^{-1})$ and introduce the notations $\widetilde{\mathbf{A}}_{\mathbf{x}}$, $\widetilde{\mathbf{A}}_{\mathbf{v}}$, and $\widehat{\mathbf{A}}_{\mathbf{v}}$ so that:
(367) $\langle\dot{\widetilde{\mathbf{x}}}\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\rangle$ $=\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\widetilde{\mathbf{x}}+\mathcal{O}(a),$
(368) $\langle\dot{\widetilde{\mathbf{v}}}\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\rangle$ $=\mu\widetilde{\mathbf{A}}_{\mathbf{v}}\widetilde{\mathbf{v}}+\mathcal{O}(a),$
(369) $\widetilde{\mathbf{x}}$ $=\mathbf{x}-\mu^{-1}\widehat{\mathbf{A}}_{\mathbf{v}}^{-1}\mathbf{v}.$
These quantities have values:
(370) $\widetilde{\mathbf{A}}_{\mathbf{x}}$ $=-\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{A}_{\mathbf{x}}+\mathcal{O}(\mu^{-2}),$
(371) $\widetilde{\mathbf{A}}_{\mathbf{v}}$ $=\mathbf{A}_{\mathbf{v}}+\mathcal{O}(\mu^{-2}),$
(372) $\widehat{\mathbf{A}}_{\mathbf{v}}$ $=\mathbf{A}_{\mathbf{v}}+\mathcal{O}(\mu^{-2}).$
Using these quantities, similarly to Eqs. (313) and (314), the covariances $\langle\widetilde{\mathbf{x}}\widetilde{\mathbf{x}}^{\mathsf{T}}\rangle$ and $\langle\widetilde{\mathbf{v}}\widetilde{\mathbf{v}}^{\mathsf{T}}\rangle$ can be calculated with corrections $\mathcal{O}(\mu)a\mathcal{O}(a,b)$ and $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)$, respectively. For covariance functions with a time-lag, we have:
(373) $\frac{\mathrm{d}}{\mathrm{d}\tau}\langle\widetilde{\mathbf{x}}(\tau)\widetilde{\mathbf{x}}(0)^{\mathsf{T}}\rangle=\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\langle\widetilde{\mathbf{x}}(\tau)\widetilde{\mathbf{x}}(0)^{\mathsf{T}}\rangle+\mathcal{O}(\mu^{0})a\mathcal{O}(a,b).$
We multiply by the integrating factor $\exp(-\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau)$ and integrate. Pretending the correction is constant, this gives:
(374) $\langle\widetilde{\mathbf{x}}(\tau)\widetilde{\mathbf{x}}(0)^{\mathsf{T}}\rangle=\exp(\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau)\langle\widetilde{\mathbf{x}}\widetilde{\mathbf{x}}^{\mathsf{T}}\rangle+\mathcal{O}(\mu)a\mathcal{O}(a,b).$
 [^21] This argument suggests that the magnitude of the correction is (at most) divided by the complex moduli of the eigenvalues of $\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}$. Actually, it is (at most) divided by the real parts of its eigenvalues.
Similarly, we have:
(375) $\langle\widetilde{\mathbf{v}}(\tau)\widetilde{\mathbf{v}}(0)^{\mathsf{T}}\rangle=\exp(\mu\widetilde{\mathbf{A}}_{\mathbf{v}}\tau)\langle\widetilde{\mathbf{v}}\widetilde{\mathbf{v}}^{\mathsf{T}}\rangle+\mathcal{O}(\mu^{0})a\mathcal{O}(a,b).$
The same correction holds for $\langle\widetilde{\mathbf{v}}(\tau)\widetilde{\mathbf{x}}(0)^{\mathsf{T}}\rangle$. Similar results hold for fourth-order moments, e.g.:
(376) $\langle\widetilde{x}^{i}\widetilde{x}^{j}\widetilde{v}^{k}\widetilde{v}^{l}\rangle$ $=\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle\langle v^{k}\widetilde{v}^{l}\rangle+\mathcal{O}(\mu^{-2})+\mathcal{O}(\mu^{0})\mathcal{O}((a,b)^{2}),$
(377) $\langle\widetilde{x}^{i}(\tau)\widetilde{x}^{j}(\tau)\widetilde{v}^{k}(\tau)\widetilde{v}^{l}(0)\rangle$ $=\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle\langle\widetilde{v}^{k}(\tau)\widetilde{v}^{l}(0)\rangle+\mathcal{O}(\mu^{-2})+\mathcal{O}(\mu^{0})\mathcal{O}((a,b)^{2}),$
etc. For $\langle\widetilde{\mathbf{x}}(\tau)\widetilde{\mathbf{v}}(0)^{\mathsf{T}}\rangle$, we have:
(378) $\frac{\mathrm{d}}{\mathrm{d}\tau}\langle\widetilde{x}^{i}(\tau)\widetilde{v}^{j}(0)\rangle$ $=\mu^{-1}(\widetilde{\mathbf{A}}_{\mathbf{x}})^{i}_{k}\langle\widetilde{x}^{k}(\tau)\widetilde{v}^{j}(0)\rangle-\mu^{-1}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[{(\widetilde{a}_{xx})^{i^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a}]\langle\widetilde{x}^{k}(\tau)\widetilde{x}^{l}(\tau)\widetilde{v}^{j}(0)\rangle$
$\quad{}-2(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[{(a_{xv})^{i^{\prime}}_{jk}+\mathcal{O}(\mu^{-1})a}]\langle\widetilde{x}^{k}(\tau)\widetilde{v}^{l}(\tau)\widetilde{v}^{j}(0)\rangle$
$\quad{}-(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[{(a_{vv})^{i^{\prime}}_{jk}+\mathcal{O}(\mu^{-1})a}]\langle\widetilde{v}^{k}(\tau)\widetilde{v}^{l}(\tau)\widetilde{v}^{j}(0)\rangle.$
Using Eq. (158), with similar arguments for the correction,
(379) $\langle\widetilde{x}^{l}(\tau)\widetilde{x}^{m}(\tau)\widetilde{v}^{j}(0)\rangle=\mathcal{O}(\mu^{-1})\mathcal{O}(a,b)+\mathcal{O}(\mu)a\mathcal{O}((a,b)^{2})=\mathcal{O}(\mu^{-1})\mathcal{O}(a,b).$
The expectations on the second and third lines of Eq. (378) consist of a portion given by Eq. (158), which has integral $\mathcal{O}(\mu^{-1})\mathcal{O}(a,b)$, plus a correction $\mathcal{O}(\mu^{0})a\mathcal{O}((a,b)^{2})$. This gives:
(380) $\langle\widetilde{\mathbf{x}}(\tau)\widetilde{\mathbf{v}}(0)^{\mathsf{T}}\rangle=\exp(\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau)\langle\widetilde{\mathbf{x}}\widetilde{\mathbf{v}}^{\mathsf{T}}\rangle+\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b).$
We will want to perform a fit:
(381) $\langle\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ $\approx(\breve{A}_{x})^{i}_{j}\widetilde{x}^{j}(0)+(\breve{A}_{v})^{i}_{j}\widetilde{v}^{j}(0)+(\breve{a}_{xx})^{i}_{jk}(\widetilde{x}^{j}(0)\widetilde{x}^{k}(0)-\langle\widetilde{x}^{j}\widetilde{x}^{k}\rangle)$
$\quad{}+2(\breve{a}_{xv})(\widetilde{x}^{j}(0)\widetilde{v}^{k}(0)-\langle\widetilde{x}^{j}\widetilde{v}^{k}\rangle)+(\breve{a}_{vv})^{i}_{jk}(\widetilde{v}^{j}(0)\widetilde{v}^{k}(0)-\langle\widetilde{v}^{j}\widetilde{v}^{k}\rangle)$
To do this, we multiply by a function $f(\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0))$ and take expectations. We choose linear and quadratic monomials for $f(\cdot)$. From the linear choices of $f(\cdot)$, we have:
(382) $\breve{\mathbf{A}}_{\mathbf{x}}$ $=\exp({\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau})-\mathds{1}+\mathcal{O}(\mu)a\mathcal{O}(a,b),$
(383) $\breve{\mathbf{A}}_{\mathbf{v}}$ $=\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b).$
For the quadratic choices of $f(\cdot)$, we have from Eq. (157):
$(\breve{a}_{xx})^{k}_{ij}$ =$=\Big{\{}{\Big{[}{(\widetilde{\mathbf{A}}_{\mathbf{x}}\otimes\mathds{1}\otimes\mathds{1}+\mathds{1}\otimes\widetilde{\mathbf{A}}_{\mathbf{x}}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes\widetilde{\mathbf{A}}_{\mathbf{x}})^{-1}}\Big{.}}\Big{.}$
.×$\Big{.}{\quad\qquad{}\times\left({\exp({\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau})\otimes\exp({\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau})\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes\exp({\mu^{-1}\widetilde{\mathbf{A}}_{\mathbf{x}}\tau})}\right)}\Big{]}^{i^{\prime}j^{\prime}k}_{ijk^{\prime}}(\widetilde{a}_{xx})^{k^{\prime}}_{i^{\prime}j^{\prime}}$
(384) .$\Big{.}{\qquad{}+\mathcal{O}(\mu^{-1})a}\Big{\}}[1+\mathcal{O}(\mu)\mathcal{O}((a,b)^{2})],$
(385) $\breve{a}_{xv}$ $=\mathcal{O}(\mu^{-1})a+\mathcal{O}(\mu)a\mathcal{O}((a,b)^{2})=\mathcal{O}(\mu^{-1})a,$
(386) $\breve{a}_{vv}$ $=\mathcal{O}(\mu^{-1})a+\mathcal{O}(\mu)a\mathcal{O}((a,b)^{2})=\mathcal{O}(\mu^{-1})a.$
We now introduce another time-lag $\tau^{\prime}>0$ and calculate using Eq. (381):
$\langle(\widetilde{x}^{i}(\tau+\tau^{\prime})-\widetilde{x}^{i}(\tau^{\prime}))(\widetilde{x}^{j}(\tau^{\prime})-\widetilde{x}^{j}(0))\rangle$
(387) $\qquad{}=\langle[(\breve{A}_{x})^{i}_{k}\widetilde{x}^{k}(\tau^{\prime})+(\breve{a}_{xx})^{i}_{kl}(\widetilde{x}^{k}(\tau^{\prime})\widetilde{x}^{l}(\tau^{\prime})-\langle\widetilde{x}^{k}\widetilde{x}^{l}\rangle)](\widetilde{x}^{j}(\tau^{\prime})-\widetilde{x}^{j}(0))\rangle+\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b),$
$\langle(\widetilde{x}^{i}(\tau+\tau^{\prime})-\widetilde{x}^{i}(\tau^{\prime}))\widetilde{x}^{j}(\tau^{\prime})(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle$
(388) $\qquad{}=\langle[(\breve{A}_{x})^{i}_{l}\widetilde{x}^{l}(\tau^{\prime})+(\breve{a}_{xx})^{i}_{lm}(\widetilde{x}^{l}(\tau^{\prime})\widetilde{x}^{m}(\tau^{\prime})-\langle\widetilde{x}^{l}\widetilde{x}^{m}\rangle)]\widetilde{x}^{j}(\tau^{\prime})(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle+\mathcal{O}(\mu^{-2})a,$
$\langle(\widetilde{x}^{i}(\tau+\tau^{\prime})-\widetilde{x}^{i}(\tau^{\prime}))(\widetilde{x}^{j}(\tau^{\prime})-\widetilde{x}^{j}(0))(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle$
(389) $\qquad{}=\langle[(\breve{A}_{x})^{i}_{l}\widetilde{x}^{l}(\tau^{\prime})+(\breve{a}_{xx})^{i}_{lm}(\widetilde{x}^{l}(\tau^{\prime})\widetilde{x}^{m}(\tau^{\prime})-\langle\widetilde{x}^{l}\widetilde{x}^{m}\rangle)](\widetilde{x}^{j}(\tau^{\prime})-\widetilde{x}^{j}(0))(\widetilde{x}^{k}(\tau^{\prime})-\widetilde{x}^{k}(0))\rangle+\mathcal{O}(\mu^{-2})a.$
For $\tau,\tau^{\prime}\lesssim\mathcal{O}(\mu)$, the l.h.s.’s of the above is compared with a magnitude of $\mathcal{O}(\mu^{-2}\tau\tau^{\prime})$. Thus, with regards to the drift, the dynamics of $\widetilde{\mathbf{x}}$ is nearly Markovian, with relative error $\mathcal{O}((\tau\tau^{\prime})^{-1})$.
Now we turn to the quadratic variation. First, we investigate:
$\frac{\mathrm{d}}{\mathrm{d}\tau}\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad{}=\langle\dot{\widetilde{x}}^{i}(\tau)(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle+\langle\dot{\widetilde{x}}^{j}(\tau)(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
@@
(390) $\qquad\quad{}+\left\langle\frac{\mathrm{d}[\widetilde{x}^{i},\widetilde{x}^{j}](\tau)}{\mathrm{d}\tau}\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\right\rangle$
$\qquad{}=\mu^{-1}(\widetilde{A}_{x})^{i}_{i^{\prime}}\langle\widetilde{x}^{i^{\prime}}(\tau)(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle+\mu^{-1}(\widetilde{A}_{x})^{j}_{j^{\prime}}\langle\widetilde{x}^{j^{\prime}}(\tau)(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-\mu^{-1}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[(\widetilde{a}_{xx})^{i^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{x}^{k}(\tau)\widetilde{x}^{l}(\tau)-\langle\widetilde{x}^{k}\widetilde{x}^{l}\rangle)(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-\mu^{-1}(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}[(\widetilde{a}_{xx})^{j^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{x}^{k}(\tau)\widetilde{x}^{l}(\tau)-\langle\widetilde{x}^{k}\widetilde{x}^{l}\rangle)(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-2(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[(a_{xv})^{i^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{x}^{k}(\tau)\widetilde{v}^{l}(\tau)-\langle\widetilde{x}^{k}\widetilde{v}^{l}\rangle)(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-2(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}[(a_{xv})^{j^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{x}^{k}(\tau)\widetilde{v}^{l}(\tau)-\langle\widetilde{x}^{k}\widetilde{v}^{l}\rangle)(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}[(a_{vv})^{i^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{v}^{k}(\tau)\widetilde{v}^{l}(\tau)-\langle\widetilde{v}^{k}\widetilde{v}^{l}\rangle)(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$
$\qquad\quad{}-(\widehat{\mathbf{A}}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}[(a_{vv})^{j^{\prime}}_{kl}+\mathcal{O}(\mu^{-1})a]\langle(\widetilde{v}^{k}(\tau)\widetilde{v}^{l}(\tau)-\langle\widetilde{v}^{k}\widetilde{v}^{l}\rangle)(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$

### v15

@@ 6.1. Preliminaries
showing that we only need the coefficient of $x^{3}$ to be $\mathcal{O}(a^{2})$ for the model to make sense, a result which could have been anticipated by symmetry. Thus, for small $a$, we can interpret the quadratic model for $\langle\dot{x}\mid x\rangle$ irrespective of any higher-order contributions. We also assume the linear model for inhomogeneous diffusion, noting similarly that in order for the diffusion function to be non-negative we can add a term $(b^{2}/4)x^{2}$, which is $\mathcal{O}(b^{2})$. Thus, to order $\mathcal{O}((a,b)^{2})$ [^15] We write $\mathcal{O}((a,b)^{2})$ to stand for $\mathcal{O}(a^{2},ab,b^{2})$., our model is:
@@ 6.2. One-dimensional case
$\qquad{}+\mathcal{O}(a^{2}){He}_{3}(x)+\mathcal{O}(a^{3}){He}_{4}(x)+\cdots,$
@@
$\qquad{}+(2a+a^{2}\mathcal{O}(a,b)){He}_{3}(x)+\mathcal{O}(a^{2}){He}_{4}(x)+\cdots,$
@@ 8.3.2. Solution of third-order quantities
$\quad{}+\mathcal{O}(\mu^{-1})(a_{vv},b_{v})+\mathcal{O}(\mu^{-2})(a_{xx},a_{xv},b_{x}),$
@@ 8.3.3. Quantitative significance
which is also a reasonable choice. Lastly, it should be noted that $\langle x^{i}v^{j}v^{k}\rangle=L(x^{i}v^{j},x^{k})/2$ is also an angular momentum, but the calculated ensemble covariance using the angular momenta is $\mathcal{O}(\mu^{-2})$ when $\mathbf{x}$ and $\mathbf{v}$ are expressed in terms of $\widetilde{\mathbf{x}}$ and $\widetilde{\mathbf{v}}$ and when Eqs. (361)–(362) are applied.
@@ 8.4. Dynamics on long time-scales
We now show that under the assumption $\left|(a,b)\right|\lesssim\mathcal{O}(\mu^{-1})$, the dynamics of $\mathbf{x}$ is Markovian on $\mathcal{O}(\mu)$ time-scales with $\mathcal{O}(\mu^{-2})$ error. We do this by solving for the generalized Koopman eigenfunctions as in subsection 6.4. First, we have $\mathbf{x}=(\widetilde{\mathbf{x}}+\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\widetilde{\mathbf{v}})(1+\mathcal{O}(\mu^{-2}))$. We will do computations with the transformed variables $\widetilde{\mathbf{x}}$, $\widetilde{\mathbf{v}}$. The generalized Koopman eigenfunctions with leading terms $\widetilde{\mathbf{x}}$ and $\widetilde{\mathbf{v}}$ are, respectively:
$\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})$ $=\widetilde{\mathbf{x}}+\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b)\widetilde{\mathbf{v}}+\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{x}^{j}-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
(366) $\quad{}+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{v}^{j}-\langle\widetilde{x}^{i}\widetilde{v}^{j}\rangle)+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{v}^{i}\widetilde{v}^{j}-\langle\widetilde{v}^{i}\widetilde{v}^{j}\rangle)+\cdots,$
$\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})$ $=\widetilde{\mathbf{v}}+a\mathcal{O}(a,b)\widetilde{\mathbf{x}}+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{x}^{j}-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
(367) $\quad{}+\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{v}^{j}-\langle\widetilde{x}^{i}\widetilde{v}^{j}\rangle)+\mathcal{O}(a)(\widetilde{v}^{i}\widetilde{v}^{j}-\langle\widetilde{v}^{i}\widetilde{v}^{j}\rangle)+\cdots,$
with dynamics:
@@
(368) $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\right\rangle$ $=-\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{A}_{\mathbf{x}}[1+\mathcal{O}(\mu^{-2})+\mathcal{O}(\mu)a\mathcal{O}(a,b)]\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}),$